OpenStudy (anonymous):

find the value of 10sqrt{2} over sqrt{5}

OpenStudy (mimi_x3):

\[\frac{10\sqrt{2}}{\sqrt{5}} =>\frac{10\sqrt{2}*\sqrt{5}}{\sqrt{5}*\sqrt{5}} = \frac{10\sqrt{10}}{5} = 2\sqrt{10} \]

OpenStudy (anonymous):

thanks :)